OpenStudy (anonymous):

Calculate the power of i. i ^48

OpenStudy (jdoe0001):

\(\bf \textit{what's }i^2?\)

OpenStudy (anonymous):

-1

OpenStudy (jdoe0001):

\(\bf 48\div 2 = 24\implies 24\cdot 2=48\\ \quad \\ \quad \\ i^{48}\implies (i^2)^{24}\implies (-1)^{24}\) what do you think?

OpenStudy (anonymous):

Ugh, I don't know. I wasnt there for the lesson. Wouldn't it be -1 still??

OpenStudy (jdoe0001):

well, -1 * -1 = ?

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

yeap if you multiply -1 an EVEN number of times, your result will be +1 if you multiply -1 an ODD number of times, then you'd get -1 notice \(\bf 48\div 2 = 24\implies 24\cdot 2=48\\ \quad \\ \quad \\ i^{48}\implies (i^2)^{24}\implies (-1)^{{\color{red}{ 24}}}\implies (-1)(-1)...{\color{red}{ 24}}\ times\)

OpenStudy (anonymous):

so it would be 1 because the 24 is even?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

okay so for any question that asks me to calculate the power of i I just divide it by 2????

OpenStudy (jdoe0001):

yeap you set it firstly as an exponential with a \(\bf i^2\) because say for example \(\bf i^6\implies i^2\cdot i^2\cdot i^2\implies -1\cdot -1\cdot -1\implies -1 \\ \quad \\ i^6\implies (i^2)^{\color{red}{3}}\) so even though the original exponent is EVEN, 6 the result is -1 if you don't set it up first as \(i^2\)

OpenStudy (anonymous):

okay. I have this problem. i^ 361 what do I do since its an odd number??

OpenStudy (jdoe0001):

let's see 361/2 = 180.5 so we can use 180 safely so 180 * 2 = 360 so \(\bf i^{361}\implies i^{360}\cdot i^1\implies (i^2)^{{\color{red}{ 180}}}\cdot i^1\implies +1\cdot i\implies i\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw